# LITP-25-17

### v1

@@ 3.2 Properties of Stripped Amplitudes
    Lorentz (little-group) covariance. This follows immediately, since $\mathcal{W}_{ij}$ is a Lorentz scalar (in dim reg., otherwise in the limit of vanishing IR regulator).
@@
    Gravity. A version of the Froissart-Martin bound [Froissart:1961ux, Martin:1962rt], originally proved for gapped theories, holds in gravitational theories  [Haring:2022cyf]  [^9] It is interesting to contrast Regge scaling of the full amplitude with the property of single partial waves, which have been shown to violate polynomial boundedness for large $\mathcal{E}$ in the context of wave scattering off black holes [Correia:2025enx].: it is not the amplitude that directly satisfies the bound, but the amplitude smeared in momentum transfer $q\in[\mathcal{E},q_{\max}]$ with a suitable function $\psi(q)$, at $\epsilon>0$. In particular, the smeared amplitude satisfies twice–subtracted dispersion relations with,
    $\oint_{|s|\to\infty}\frac{ds}{s}\int_{\mathcal{E}}^{q_{\max}}dq\,q^{2\epsilon}\,\psi(q)\,\frac{\mathcal{M}^{\epsilon}(s,-q^{2})}{s^{2}}\;\longrightarrow\;0\,$, (3.9)
    The proof relies on the eikonal expansion of the amplitude at sufficiently large and real values of $s$ [Haring:2022cyf]. Eq. (3.9) carries over to the stripped amplitude $\mathcal{M}_{\mathcal{E}}$, defined in the limit $\epsilon\to 0$, because the eikonal regime can be extended outside the real axis—for large complex values of $s$:
@@
    Indeed, the diagrammatic proof of leading eikonal amplitude, which selects ladders and cross-ladder diagrams controlled by $G$, does not rely on the reality of $s$. Under such an extension, the (smeared) stripped amplitude has the asymptotic form,
@@
    where the divergent $\sim 1/\epsilon$ part has canceled between the soft factor (3.4) and the amplitude (3.10). Performing the integral explicitly over the big circle at infinity, we find that the Regge bound holds for the (smeared) stripped amplitude as well:
    $\oint_{|s|\to\infty}\frac{ds}{s}\int_{\mathcal{E}}^{q_{\max}}dq\,\psi(q)\,\frac{\mathcal{M}_{\mathcal{E}}(s,-q^{2})}{s^{2}}\;\longrightarrow\;0\,.$ (3.12)
@@
    $\lim_{|s|\to+\infty}\mathcal{M}^{\epsilon}/s^{2}\to 0\,,\qquad t\ \text{fixed}\,$, (3.13)
@@
    $\lim_{|s|\to+\infty}\mathcal{M}^{\epsilon}_{\mathcal{E}}/s^{2}\to 0\,,\qquad t\ \text{fixed}\,.$ (3.14)
@@
    As for the Regge scaling of $\mathcal{M}_{\mathcal{E}}=\lim_{\epsilon\to 0^{+}}\mathcal{M}_{\mathcal{E}}^{\epsilon}$, that is whether the $\epsilon\to 0$ and $|s|\to\infty$ limits commute, we can not prove it in full generality. However, as the only gapless particles in this theory are spin-$1$, we expect no violation of the Froissart bound. Indeed, the leading contribution of gapless modes to the stripped amplitude in the eikonal regime $t\ll|s|$, with QED-only insertions, scales as $\alpha s/t$, which is bounded by $s^{2}$. Supported by this intuition, we assume,
    $\oint_{|s|\to\infty}\frac{ds}{s}\int_{\mathcal{E}}^{q_{\max}}dq\psi(q)\frac{\mathcal{M}_{\mathcal{E}}(s,-q^{2})}{s^{2}}\to 0\,$, (3.15)
    which allows us to derive twice–subtracted dispersion relations for $\mathcal{M}_{\mathcal{E}}$ in the next sections.
@@ 4 Unitarity of Stripped and Hard Amplitudes
The properties of the stripped amplitudes discussed above follow directly from those of the full amplitude and from the structure of the exponential factor. We now need to establish an appropriate notion of unitarity for $\mathcal{M}_{\mathcal{E}}$. Stripped amplitudes depend effectively on two couplings (in addition to other short-range interactions): $\alpha$ and $\alpha_{\mathcal{E}}=\alpha\log(M/\mathcal{E})$ with electromagnetism (or $GE^{2}$ and $G_{\mathcal{E}}=GM^{2}\log(M/\mathcal{E})$ in gravity) – $M$ a hard scale. We are interested in the regime of (1.1), $4\pi\gtrsim\alpha_{\mathcal{E}}\gg\alpha$ (or $4\pi\gtrsim G_{\mathcal{E}}\gg GM^{2}$), and aim to control unitarity to all orders in $\alpha_{\mathcal{E}}$ (or $G_{\mathcal{E}}$) and any other short–distance coupling, while perturbatively in $\alpha$ (or $G$):
@@ 4.1 Unitarity of Hard Amplitudes
Once a Hermitian Hamiltonian acting on a positive-definite space is constructed, the generalized optical theorem follows immediately from $S^{\mathrm{hard}\,\dagger}_{\mathcal{E}}S^{\mathrm{hard}}_{\mathcal{E}}=\mathbb{I}$ and $S^{\mathrm{hard}}_{\mathcal{E}}=\mathbb{I}+iT$  [^13] The hard hamiltonian corresponds to short–range interactions with partial waves decaying as $\ell^{-3/2}$, allowing the splitting between identity and interactions., where the identity is restricted to hard photons and gravitons, but otherwise acts as usual on the rest of the states. In particular, for a $2\to 2$ hard amplitude, the generalized optical theorem is
@@
where the sum over intermediate states $X$ excludes soft photons and gravitons, and we introduced a generalized imaginary part [Caron-Huot:2022ugt],
@@
The previous arguments are at the operator level, relying on $H^{\mathrm{hard}}_{\mathcal{E}}$. They can as well be recasted diagrammatically, in a manifest form, via cutting rules. The derivation from the largest-time equation [Veltman:1963th] remains indeed valid in the presence of $\mathcal{E}$; one can e.g. follow closely the treatment of [Anselmi:2016fid] with an appropriate gauge fixing. The only modification in the cutting rules is the replacement of the propagator $\Delta(k)$ by $\theta(|\boldsymbol{k}|-\mathcal{E})\Delta(k)$, which does not alter causality, energy scaling, or the locality of interactions.
@@ 4.2 Functionals: Unitarity and 1/t1/t–pole
At tree-level – order $O(\alpha)$ – Eq. (4.10) is manifestly valid, since the $1/t$ contribution to $\mathcal{M}_{\mathcal{E}}$ originates from the hard photon exchange. At loop–level, the full amplitude receives insertions of hard, leading soft, and subleading soft photons (or gravitons). The subleading soft insertions clearly provide further suppression by powers of $\alpha$ or $G$, and drop out after taking the scaling limit, whereas the leading soft modes are factored out by the Weinberg exponential in the definition of $\mathcal{M}_{\mathcal{E}}$. Therefore, we conclude that any left–over correction relevant for functional unitarity of $\mathcal{M}_{\mathcal{E}}$ is controlled by $\mathcal{M}^{\mathrm{hard}}_{\mathcal{E}}$.
@@ 4.3 Collinear enhancements
At sufficiently high energies, however, collinear effects may re–emerge in the form of finite but large logarithms of the type $\alpha\log(E/m)$ in QED. There are no collinear divergences instead in gravity, see e.g. [Beneke:2022pue, Akhoury:2011kq]. One may thus wonder whether the $O(\alpha)$ contributions in QED that we have systematically discarded in order to prove unitarity could, after being enhanced by such collinear logarithms, compete with the $\alpha_{\mathcal{E}}$ terms we have instead retained.
@@ 4.4 Connection to Faddeev-Kulish Amplitudes
In the scaling limit (1.1) all IR-finite amplitudes $\mathcal{M}^{FK}_{\mathcal{E}}$ defined à la Faddeev-Kulish [Kulish:1970ut, Kapec:2017tkm, Hannesdottir:2019opa] match the stripped amplitude $\mathcal{M}_{\mathcal{E}}$ by construction. Indeed the leading-order effect of the modified asymptotic dynamics of Faddeev–Kulish is exactly to remove the virtual [^15] As for additional real soft photons, their contribution vanishes in FK amplitudes in the limit (1.1). soft photon factors, namely
@@
This can be shown by repeating Weinberg’s computation and taking into account soft photons from the dressing operators. A scale $\mathcal{E}$ is inevitably introduced in the construction of IR-finite amplitudes, independently of the specific modeling of the asymptotic dynamics, either as an hard cut-off or as a renormalization-like scale [Hannesdottir:2019opa]. For instance, a possible choice—among infinitely many— for the asymptotic hamiltonian is the soft–hamiltonian itself, $H-H^{\mathrm{hard}}_{\mathcal{E}}$.
@@
In this sense, the leading contribution of $\mathcal{M}_{\mathcal{E}}$, isolated through the limit (1.1), captures the universal leading contribution to the scattering. For generic exclusive quantities, even when IR finite, corrections will instead be sensitive to the very details of the asymptotic dynamics, as it is evident from the ambiguities in Faddeev-Kulish-like amplitudes. This sensitivity disappears when considering sufficiently inclusive observables [Forde:2003jt, Hannesdottir:2019opa, Agarwal:2021ais] or energy–energy correlators [Moult:2025nhu].
@@ 5 Dispersion Relations
On the other hand, for amplitudes that are $t\leftrightarrow u$ symmetric (like $\pi^{+}\pi^{+}\to\pi^{+}\pi^{+}$), or stus-t-u symmetric (like $\pi^{0}\pi^{0}\to\pi^{0}\pi^{0}$), the definition (5.1) might equally well describe crossing symmetric dispersion relations [Bellazzini:2025shd, Auberson:1972prg, Sinha:2020win, Li:2023qzs]. In the former case, for instance, we identify ${\mathcal{M}}_{\mathcal{E}}(s,q)$ with $\mathcal{M}_{\mathcal{E}}(s,t=t_{+}(s,q))$, where,
@@
UV/IR relations follow from identifying a scale $M$ that separates the IR $|s|\leq M^{2}$, where the amplitude is computable in an EFT in terms of its Wilson coefficients and coupling constants, and the UV $|s|>M^{2}$ with unknown unitary and causal physics. [^16] The UV/IR terminology for arcs can be slightly misleading, since $|t|<M^{2}$, meaning that the momentum exchange always lies in the IR. Only $s$ probes UV regions. For instance, in gravity, at $s\gg 1/G$ and fixed $t$, the amplitude is IR-calculable through the eikonal expansion [Haring:2022cyf, Bellazzini:2022wzv]. The contour $s\in C$ lies entirely in the IR. We can then provide a UV representation of arcs by deforming $C$ around the branch cuts associated with the physical regions of the amplitude in the relevant channels $k=s,u,t$, into a large circle at infinity. The latter vanishes thanks to Regge scaling (3.15), and we obtain, using crossing symmetry,
@@
$\mathcal{A}_{n}[\psi,\mathcal{M}_{\mathcal{E}}]=\int_{M^{2}}^{+\infty}\frac{ds}{\pi}\int_{\mathcal{E}}^{q_{\max}}dq\psi(q)\left[\frac{\overline{\text{Im}}\,\mathcal{M}_{\mathcal{E}}^{s}(s,-q^{2})}{(s-s_{0})^{3+n}}+\frac{(-1)^{n}\overline{\text{Im}}\,{\mathcal{M}}^{u}_{\mathcal{E}}(s,-q^{2})}{(s+t-4m^{2}+s_{0})^{3+n}}\right]\,$, (5.5)
@@
$\mathcal{A}_{n}[\psi,\mathcal{M}_{\mathcal{E}}]=\lim_{\begin{subarray}{c}\alpha\to 0\\ \mathcal{E}\to 0\\ \alpha\log(M/\mathcal{E})\to\alpha_{\mathcal{E}}\end{subarray}}\sum_{\mathrm{cuts}\,k}\int_{M_{k}^{2}}^{+\infty}\frac{ds}{\pi}\int_{\mathcal{E}}^{q_{\max}}dq\,\psi(q){(\overline{\text{Im}}_{k}[\mathcal{M}^{k,\text{hard}}_{\mathcal{E}}}]\mathcal{K}^{k}_{n})(s,q)\,$, (5.7)
@@
Following [Caron-Huot:2021rmr], bounds on EFT couplings can thus be derived by identifying functionals $\psi(q)$ such that we get positive contributions,
@@
These bounds are conceptually similar to those obtained in large-$N$ theories [Albert:2022oes, Fernandez:2022kzi, Berman:2023jys], which hold only up to $O(1/N^{2})$ corrections, analogous to our $O(\alpha)$ corrections. The situation here is perhaps more satisfactory, since increasing the ratio $\alpha_{\mathcal{E}}/\alpha$ does not require changing the theory itself, $N$, but rather adjusting the experimental resolution $\mathcal{E}$ that probes the same underlying dynamics.
Before ending this section, it is instructive to contrast quantitatively the obstruction to exact positivity bounds—the amount of negativity—found in earlier works [Chang:2025cxc, Caron-Huot:2021rmr, Henriksson:2022oeu, Caron-Huot:2022ugt] with our finding, where the negativity is finite and of order $O(GM^{2})$ or $O(\alpha)$. In our language, the gravitational negativity of arcs identified in those analyses, may be quantified by an integral (to be further integrated in $s$) analogous to (5.10),
@@
where $\lambda$ is an IR regulator that is eventually sent to zero. Positive functionals, in the sense of (5.13), require that the integral extends down to $q=0$. In this expression, the function $F$ is such that the limits $G\!\to\!0$ and $\log(\mathcal{E}/\lambda)\!\to\!\infty$ do not commute: one ordering is trivial, while the opposite yields a finite, calculable amount of negativity after resumming large log’s [Chang:2025cxc], but leads to a puzzling non-decoupling effect of gravity, a point to which we return in the conclusions. In our framework, instead, Eq. (5.10) has the same form as (5.18), but with $\lambda\to\mathcal{E}$, so the integral is never logarithmically enhanced—an additional benefit of working directly with IR–finite amplitudes. As a result, the correction from gravity is small and connects smoothly to the unperturbed theory. The same conclusion holds for electromagnetism.

### v2

@@ 3.2 Properties of Stripped Amplitudes
    Lorentz (little-group) covariance. This follows immediately, since $\mathcal{W}_{ij}$ is a Lorentz scalar. In particular, the external momenta enter in $\mathcal{W}_{ij}$ in the Lorentz invariant combination $s_{ij}$.
@@
    Gravity. A version of the Froissart-Martin bound [81, 115], originally proved for gapped theories, holds in gravitational theories  [93]  [^10] It is interesting to contrast Regge scaling of the full amplitude with the property of single partial waves, which have been shown to violate polynomial boundedness for large $\mathcal{E}$ in the context of wave scattering off black holes [59].: it is not the amplitude that directly satisfies the bound, but the amplitude smeared in momentum transfer $q\in[\mathcal{E},q_{\max}]$ with a suitable function $\psi(q)$, at $\epsilon>0$. In particular, the smeared amplitude satisfies twice–subtracted dispersion relations, for $n\geq 0$
    $\oint_{|s|\to\infty}\frac{ds}{s}\int_{\mathcal{E}}^{q_{\max}}dq\,q^{2\epsilon}\,\psi(q)\,\frac{\mathcal{M}^{\epsilon}(s,-q^{2})}{s^{2+n}}\;\longrightarrow\;0\,.$ (3.9)
    The proof relies on the eikonal expansion of the amplitude at sufficiently large and real values of $s$ [93], see also [46]. A result analogous to Eq. (3.9) can be obtained for the stripped amplitude $\mathcal{M}_{\mathcal{E}}$, defined in the limit $\epsilon\to 0$, even without smearing in $q$, because the eikonal regime can be extended outside the real axis—for large complex values of $s$:
@@
    Indeed, the diagrammatic proof of leading eikonal amplitude, which selects ladders and cross-ladder diagrams controlled by $G$, does not rely on the reality of $s$. Under such an extension, the stripped amplitude has the asymptotic form,
@@
    where the divergent $\sim 1/\epsilon$ part has canceled between the soft factor (3.4) and the amplitude (3.10). Performing the ss–integration over the large circle at infinity explicitly, as in Eq. (3.9), we find that it vanishes for odd $n$, while for even (and non-negative) $n$ it is proportional to $\pi-2\mathrm{Si}(G|s|\log(q^{2}/\mathcal{E}^{2}))+O(1/|s|)$. In fact the integral over the large circle decays as $1/|s|^{n+1}$. We conclude that an average Regge bound also holds for the stripped amplitude: for $n\geq 0$ it yields
    $\oint_{|s|\to\infty}\frac{ds}{s}\int_{\mathcal{E}}^{q_{\max}}dq\,\psi(q)\,\frac{\mathcal{M}_{\mathcal{E}}(s,-q^{2})}{s^{2+n}}\;\longrightarrow\;0\,.$ (3.12)
    The same conclusion can be reached following similar steps to [93, 46] where one makes use of the smearing wavepacket $\psi(q)$ and of the eikonal formula for real kinematics. In our case, however, unitarity of the stripped amplitude is only approximate, see Section 4, and therefore the Regge scaling is recovered—with this method, as opposed to the previous one—only in the scaling limit (1.1).
@@
    $\lim_{|s|\to+\infty}\mathcal{M}^{\epsilon}/s^{2}=0\,,\qquad t\ \text{fixed}\,$, (3.13)
@@
    $\lim_{|s|\to+\infty}\mathcal{M}^{\epsilon}_{\mathcal{E}}/s^{2}=0\,,\qquad t\ \text{fixed}\,.$ (3.14)
@@
    As for the Regge scaling of $\mathcal{M}_{\mathcal{E}}=\lim_{\epsilon\to 0^{+}}\mathcal{M}_{\mathcal{E}}^{\epsilon}$, that is whether the $\epsilon\to 0$ and $|s|\to\infty$ limits commute, we can not prove it in full generality. However, as the only gapless particles in this theory are spin-$1$, we expect no violation of the Froissart bound. Indeed, the leading contribution of gapless modes to the stripped amplitude in the eikonal regime $t\ll|s|$, with QED-only insertions, scales as $\alpha s/t$, which is bounded by $s^{2}$. Supported by this intuition, we assume that for $n\geq 0$
    $\oint_{|s|\to\infty}\frac{ds}{s}\int_{\mathcal{E}}^{q_{\max}}dq\psi(q)\frac{\mathcal{M}_{\mathcal{E}}(s,-q^{2})}{s^{2+n}}\to 0\,$, (3.15)
    which allows us to derive at least twice–subtracted dispersion relations for $\mathcal{M}_{\mathcal{E}}$ in the next sections.
    Alternatively, as for the gravitational case, an approximate version of this conclusion can be reached following the steps of [93, 46], where the smearing wavepacket $\psi(q)$ allows to control the large partial–waves contribution.
@@ 4 Unitarity of Stripped and Hard Amplitudes
The properties of the stripped amplitudes discussed above follow directly from those of the full amplitude and from the structure of the exponential factor. We now need to establish an appropriate notion of unitarity for $\mathcal{M}_{\mathcal{E}}$. Stripped amplitudes depend effectively on two couplings (in addition to other short-range interactions): $\alpha$ and $\alpha_{\mathcal{E}}=\alpha\log(M/\mathcal{E})$ with electromagnetism (or $GM^{2}$ and $G_{\mathcal{E}}=GM^{2}\log(M/\mathcal{E})$ in gravity) – $M$ a hard scale. We are interested in the regime of (1.1), $4\pi\gtrsim\alpha_{\mathcal{E}}\gg\alpha$ (or $4\pi\gtrsim G_{\mathcal{E}}\gg GM^{2}$), and aim to control unitarity to all orders in $\alpha_{\mathcal{E}}$ (or $G_{\mathcal{E}}$) and any other short–distance coupling, while perturbatively in $\alpha$ (or $G$):
@@ 4.1 Unitarity of Hard Amplitudes
Once a Hermitian Hamiltonian acting on a positive-definite space is constructed, the generalized optical theorem follows immediately from $S^{\mathrm{hard}\,\dagger}_{\mathcal{E}}S^{\mathrm{hard}}_{\mathcal{E}}=\mathbb{I}$ and $S^{\mathrm{hard}}_{\mathcal{E}}=\mathbb{I}+iT$  [^14] The hard hamiltonian corresponds to short–range interactions with partial waves decaying as $\ell^{-3/2}$ times an oscillatory factor that helps convergence, allowing the splitting between identity and interactions., where the identity is restricted to hard photons and gravitons, but otherwise acts as usual on the rest of the states. In particular, for a $2\to 2$ hard amplitude, the generalized optical theorem is
@@
where the sum over intermediate states $X$ excludes soft photons and gravitons, and we introduced a generalized imaginary part [46],
@@
The previous arguments are at the operator level, relying on $H^{\mathrm{hard}}_{\mathcal{E}}$. They can as well be recasted diagrammatically, in a manifest form, via cutting rules. The derivation from the largest-time equation [140] remains indeed valid in the presence of $\mathcal{E}$; one can e.g. follow closely the treatment of [10] with an appropriate gauge fixing. The only modification in the cutting rules is the replacement of the propagator $\Delta(k)$ by $\theta(|\boldsymbol{k}|-\mathcal{E})\Delta(k)$, which does not alter causality, energy scaling, or the locality of interactions.
@@ 4.2 Functionals: Unitarity and 1/t1/t–pole
At tree-level —order $O(\alpha)$ or order O(G)O(G)—Eq. (4.10) is manifestly valid, since the $1/t$ contribution to $\mathcal{M}_{\mathcal{E}}$ originates from the hard photon (graviton) exchange. At loop–level, the full amplitude receives insertions of hard, leading soft, and subleading soft photons (or gravitons). The subleading soft insertions clearly provide further suppression by powers of $\alpha$ or $G$, and drop out after taking the scaling limit, whereas the leading soft modes are factored out by the Weinberg exponential in the definition of $\mathcal{M}_{\mathcal{E}}$. Therefore, we conclude that any left–over correction relevant for functional unitarity of $\mathcal{M}_{\mathcal{E}}$ is controlled by $\mathcal{M}^{\mathrm{hard}}_{\mathcal{E}}$.
@@
Before ending this subsection it is perhaps worth stressing that other approximation schemes may apply under extra assumptions. An obvious example is when the stripped amplitude $\mathcal{M}_{\mathcal{E}}$ is assumed to be well approximated by a meromorphic, tree–level amplitude even in the UV. This is similar to the limit that is considered in the bootstrap program of large–N gauge theories [6, 79, 34] or in certain tree–level stringy UV completion of Einstein–gravity, e.g. [44]. In these special cases, even the orders $O(\alpha)$ and $O(G)$ are under control and manifestly unitary, $\mathcal{M}_{\mathcal{E}}\big|_{\mathrm{tree}}=\mathcal{M}^{\epsilon}\big|_{\mathrm{tree}}=\mathcal{M}^{\mathrm{hard}}_{\mathcal{E}}\big|_{\mathrm{tree}}$ for $\mathcal{E}/M\ll 1$, giving rise to $\int_{\mathcal{E}}^{q_{\max}}\!\!dq~\psi(q)\,\overline{\text{Im}}\,\mathcal{M}_{\mathcal{E}}(s,-q^{2})\succeq 0$ up to power corrections $O(\mathcal{E}/M)$ and IR–finite loops. The accuracy is to order $O(G,\alpha)$ included, with error starting at $O(G^{2},\alpha^{2},g_{*}^{2}\alpha,g_{*}^{2}G,\alpha\alpha_{\mathcal{E}},\ldots)$. Importantly, as opposed to earliest studies involving $\mathcal{M}^{\epsilon}\big|_{\mathrm{tree}}$, the corrections to $\mathcal{M}_{\mathcal{E}}\big|_{\mathrm{tree}}$ are IR–finite and can be small in these class of theories.
@@ 4.3 Collinear enhancements
At sufficiently high energies, however, collinear effects may re–emerge in the form of finite but large logarithms of the type $\alpha\log(E/m)$ in QED. There are no collinear divergences instead in gravity, see e.g. [31, 3]. One may thus wonder whether the $O(\alpha)$ contributions in QED that we have systematically discarded in order to prove unitarity could, after being enhanced by such collinear logarithms, compete with the $\alpha_{\mathcal{E}}$ terms we have instead retained.
@@ 4.4 Connection to Faddeev-Kulish Amplitudes
In the scaling limit (1.1) all IR-finite amplitudes $\mathcal{M}^{FK}_{\mathcal{E}}$ defined à la Faddeev-Kulish [109, 104, 91] match the stripped amplitude $\mathcal{M}_{\mathcal{E}}$ by construction. Indeed the leading-order effect of the modified asymptotic dynamics of Faddeev–Kulish is exactly to remove the virtual [^16] As for additional real soft photons, their contribution vanishes in FK amplitudes in the limit (1.1). soft photon factors, namely
@@
This can be shown by repeating Weinberg’s computation and taking into account soft photons from the dressing operators. A scale $\mathcal{E}$ is inevitably introduced in the construction of IR-finite amplitudes, independently of the specific modeling of the asymptotic dynamics, either as an hard cut-off or as a renormalization-like scale [91]. For instance, a possible choice—among infinitely many— for the asymptotic hamiltonian is the soft–hamiltonian itself, $H-H^{\mathrm{hard}}_{\mathcal{E}}$.
@@
In this sense, the leading contribution of $\mathcal{M}_{\mathcal{E}}$, isolated through the limit (1.1), captures the universal leading contribution to the scattering. For generic exclusive quantities, even when IR finite, corrections will instead be sensitive to the very details of the asymptotic dynamics, as it is evident from the ambiguities in Faddeev-Kulish-like amplitudes. This sensitivity disappears when considering sufficiently inclusive observables [^17] In fact, as discussed in [45], interference effects are IR sensitive even for inclusive observables, leading to decoherence unless the states are properly dressed. [80, 91, 2] or energy–energy correlators [118].
@@ 5 Dispersion Relations
On the other hand, for amplitudes that are $t\leftrightarrow u$ symmetric (like $\pi^{+}\pi^{+}\to\pi^{+}\pi^{+}$), or stus-t-u symmetric (like $\pi^{0}\pi^{0}\to\pi^{0}\pi^{0}$), the definition (5.1) might equally well describe crossing symmetric dispersion relations [25, 15, 133, 111]. In the former case, for instance, we identify ${\mathcal{M}}_{\mathcal{E}}(s,q)$ with $\mathcal{M}_{\mathcal{E}}(s,t=t_{+}(s,q))$, where,
@@
UV/IR relations follow from identifying a scale $M$ that separates the IR $|s|\leq M^{2}$, where the amplitude is computable in an EFT in terms of its Wilson coefficients and coupling constants, and the UV $|s|>M^{2}$ with unknown unitary and causal physics. [^18] The UV/IR terminology for arcs can be slightly misleading, since $|t|<M^{2}$, meaning that the momentum exchange always lies in the IR. Only $s$ probes UV regions. For instance, in gravity, at $s\gg 1/G$ and fixed $t$, the amplitude is IR-calculable through the eikonal expansion [93, 22]. The contour $s\in C$ lies entirely in the IR. We can then provide a UV representation of arcs by deforming $C$ around the branch cuts associated with the physical regions of the amplitude in the relevant channels $k=s,u,t$, into a large circle at infinity. The latter vanishes thanks to Regge scaling (3.15), and we obtain, using crossing symmetry,
@@
$\mathcal{A}_{n}[\psi,\mathcal{M}_{\mathcal{E}}]=\int_{M^{2}}^{+\infty}\frac{ds}{\pi}\int_{\mathcal{E}}^{q_{\max}}dq\psi(q)\left[\frac{\overline{\text{Im}}\,\mathcal{M}_{\mathcal{E}}^{s}(s,-q^{2})}{(s-s_{0})^{3+n}}+\frac{(-1)^{n}\overline{\text{Im}}\,{\mathcal{M}}^{u}_{\mathcal{E}}(s,-q^{2})}{(s-q^{2}-4m^{2}+s_{0})^{3+n}}\right]\,$, (5.5)
@@
$\mathcal{A}^{(0)}_{n}[\psi,\mathcal{M}_{\mathcal{E}}]=\lim_{\begin{subarray}{c}\alpha\to 0\\ \mathcal{E}\to 0\\ \alpha\log(M/\mathcal{E})\to\alpha_{\mathcal{E}}\end{subarray}}\sum_{\mathrm{cuts}\,k}\int_{M_{k}^{2}}^{+\infty}\frac{ds}{\pi}\int_{\mathcal{E}}^{q_{\max}}dq\,\psi(q){(\overline{\text{Im}}_{k}[\mathcal{M}^{k,\text{hard}}_{\mathcal{E}}}]\mathcal{K}^{k}_{n})(s,q)\,$, (5.7)
@@
Following [49], bounds on EFT couplings can thus be derived by identifying functionals $\psi(q)$ such that we get positive contributions,
@@
These bounds are conceptually similar to those obtained in large-$N$ theories [6, 79, 34], which hold only up to $O(1/N^{2})$ corrections, analogous to our $O(\alpha)$ corrections. The situation here is perhaps more satisfactory, since increasing the ratio $\alpha_{\mathcal{E}}/\alpha$ does not require changing the theory itself, $N$, but rather adjusting the experimental resolution $\mathcal{E}$ that probes the same underlying dynamics.
Before ending this section, it is instructive to contrast quantitatively the obstruction to exact positivity bounds—the amount of negativity—found in earlier works [53, 49, 98, 46] with our finding, where the negativity is finite and of order $O(GM^{2})$ or $O(\alpha)$. In our language, the gravitational negativity of arcs identified in those analyses, may be quantified by an integral (to be further integrated in $s$) analogous to (5.10),
@@
where $\lambda$ is an IR regulator that is eventually sent to zero. Positive functionals, in the sense of (5.13), require that the integral extends down to $q=0$. In this expression, the function $F$ is such that the limits $G\!\to\!0$ and $\log(\mathcal{E}/\lambda)\!\to\!\infty$ do not commute: one ordering is trivial, while the opposite yields a finite, calculable amount of negativity after resumming large log’s [53], but leads to a puzzling non-decoupling effect of gravity, a point to which we return in the conclusions. Setting $\mathcal{E}\simeq\lambda$ is of no help, because while the negativity is no longer harmful, the smearing of the IR side of the dispersion reintroduce the IR sensitivity of the form $\overline{F}(GM^{2}\log M/\lambda)$ and the limits $G\to 0$ and $\lambda\to 0$ again do not commute.
In our framework, instead, Eq. (5.10) has the same form as (5.18), but with the effective replacement $\lambda\to\mathcal{E}$, so the integral is never logarithmically enhanced—an additional benefit of working directly with IR–finite amplitudes. In practice, the $GM^{2}\log M/\mathcal{E}$ can be chosen to be small for a finite detector resolution $\mathcal{E}$. As a result, the correction from gravity connects smoothly to the unperturbed theory. The same conclusion holds for electromagnetism.